# Limit Theorems For Non-Hermitian Ensembles

### v3

@@ A.3. Proof of Theorem 10
$e^{-\eta\alpha N\left(\frac{k}{\alpha N}-1\right)-\alpha N\ln(1-\eta)}$ $\geq e^{(\alpha N+k)(\eta-\ln(1-\eta))}$
$=e^{-\eta\alpha N(\frac{k}{\alpha N}-1)-\left(\alpha N+k\right)\ln(1-\eta)}>1$
And with $0<\eta<1$ and $\forall Y\in\mathbb{R}$
$e^{Y-\eta\sqrt{\alpha N}\phi_{N}^{(R)}(x)}<e^{Y}\Rightarrow 1-e^{Y-\eta\sqrt{\alpha N}\phi_{N}^{(R)}(x)}>1-e^{Y}$
Now, setting $\eta=1-\frac{1}{\frac{k}{\alpha N}-1}$ which does maximise the remaining exponent (similarly presented in [20]) and applying the the quantile formula of the exponential distribution, this implies
$\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
$\geq\prod_{k=\alpha N\delta_{N}}^{N}\left(1-e^{-\alpha N\left[\eta\left(\frac{k}{\alpha N}-1\right)+\ln(1-\eta)\right]}\right)$
$\geq\prod_{k=\alpha N\delta_{N}}^{N}\left(1-e^{-\alpha N\left[\frac{k}{\alpha N}-\ln\left(\frac{k}{\alpha N}-1\right)\right]}\right)$
$\geq\left(1-e^{-\alpha N\delta_{N}^{2}}\right)^{N}$
Finally,
$\prod_{k=\alpha N\delta_{N}}^{N}p_{k}\geq\left(1-e^{-\alpha N\delta_{N}^{2})}\right)^{N}$
The parameter $\delta_{N}$ is chosen such that $\left(1-e^{-\alpha N\delta_{N}^{2})}\right)^{N}=\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$, i.e., $\delta_{N}=\sqrt{\frac{2\log{N}}{\alpha N}}$.
The partial product $\prod_{k=1}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$ is bounded. It is established from the following that the limit of the partial product $\prod_{k=1}^{\alpha N\delta_{N}}h_{k}$ is equal to one as $N$ goes infinity.
More precisely,
$\prod_{k=1}^{\alpha N\delta_{N}}h_{k}\prod_{k=\alpha N\delta_{N}}^{N}h_{k}\leq\prod_{k=1}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]\leq\prod_{k=1}^{\alpha N\delta_{N}}h_{k}$
Lower bound of the partial product $\prod_{k=\alpha N\delta_{N}}^{N}h_{k}$.
For large $N$, as the random variable $\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)$ follows a standard Gaussian distribution, and taking into account a similar remark as in the paper of B. Rider [20] saying that the upper limit of integration can be extended to $\infty$, the probability
$P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\approx 1$
Now, considering the probability
$P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)$
$=1-P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)>-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)$
$=1-P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)<\phi_{N}^{(r)}(y)+\frac{k}{\sqrt{\alpha N}}\right)<\frac{1}{N^{2}}$
where the random variable $\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)$ follows a standard Gaussian distribution.
This implies that
$P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)>-\frac{1}{N^{2}}$
Consequently,
$1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}>1-\frac{1}{N^{2}}$
Thus,
$\prod_{k=\alpha N\delta_{N}}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$
$>\prod_{k=\alpha N\delta_{N}}^{N}\left(1-\frac{1}{N^{2}}\right)>\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$
The lower bound is $1-O\left(\frac{1}{N}\right)$ and goes to $1$ as $N$ goes to infinity.
@@
$\log\prod_{k=1}^{\alpha N\delta_{N}}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+\alpha N}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
$=\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{k+\alpha N}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
@@
Furthermore, using the classical Edgeworth expansion
$\log P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
$=\log\left[\int_{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}^{\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{\alpha N-k}}\sup_{|c|\leq Y}(\phi_{N}^{(R)}(c))^{2}e^{-\frac{(\phi_{N}^{(R)}(c))^{2}}{2}}\right)$
$+O\left(\frac{1}{\alpha N-k}\right)+O\left(\phi_{N}^{(r)}(y)(\alpha N-k)^{-3/2}\right)+O\left(e^{-\frac{(\phi_{N}^{(r)}(y))^{2}}{2}}\right)$
@@
where, as in [20], for $x$ restricted as in $|x|\leq Y$ for some large positive $Y$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq Y}\phi_{N}(x)$.
The leading order sum as the Riemann integral
$\sum_{k=1}^{\sqrt{2N\log N}}\log\left[\int_{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}^{\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$ $\simeq\sqrt{\alpha N}\int_{\phi_{N}^{(R)}(x)}^{+\infty}\log\left[\int_{-\infty}^{u}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du$
$+\sqrt{\alpha N}\int_{\phi_{N}^{(r)}(y)}^{+\infty}\log\left[\int_{-u}^{+\infty}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du+E_{N}$
@@
$E_{N}=O\left(\left(\sqrt{\frac{\log N}{\alpha N}}\right)\vee\left(\sqrt{\log{N}}\sup_{|c|\leq Y}(\phi_{N}^{(R)}(c))^{2}e^{-\frac{(\phi_{N}^{(R)}(c))^{2}}{2}}\right)\right)$
Following the framework presented in [20] and detailed in Appendices A.1 and A.2,
$\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(R)}(x)}^{+\infty}\log\left[\int_{-\infty}^{u}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du=\log F_{X}^{(\infty)}(x)=-\exp(-x)$
and
$\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(r)}(y)}^{+\infty}\log\left[\int_{-u}^{+\infty}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du=\log\left(1-F_{Y}^{(\infty)}(y)\right)=-\exp(y)$
@@
$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right]$
@@
This proves the independence of the scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$ for $N\times N$ matrices from the complex induced Ginibre ensemble with proportional rectangularity index as $N$ goes to infinity.

### v4

@@ A.3. Proof of Theorem 10
$\prod_{k=\alpha N\delta_{N}}^{N}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}\left[P\left(\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)-P\left(\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)\right]$
$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\geq\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}\right)-P\left(\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)\right]$
$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\geq\sqrt{\alpha N}\Phi_{N}^{(R)}(x)-k\right)-P\left(\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)\right]$
$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{k+L}Z_{j}\geq\sqrt{\alpha N}\Phi_{N}^{(R)}(x)+\alpha N\right)-P\left(\frac{\sum_{j=1}^{k+L}Z_{j}}{\sqrt{\alpha N+k}}\leq-\left(\sqrt{\frac{\alpha N}{\alpha N+k}}\right)\phi_{N}^{(r)}(y)+\frac{k}{\alpha N+k}\right)\right]$
$>\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{k+L}Z_{j}\geq\sqrt{\alpha N}\Phi_{N}^{(R)}(x)+\alpha N\right)-P\left(\frac{\sum_{j=1}^{k+L}Z_{j}}{\sqrt{\alpha N+k}}\leq\frac{k}{\alpha N+k}\right)\right]$
$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{k+L}Z_{j}\geq\sqrt{\alpha N}\Phi_{N}^{(R)}(x)+\alpha N\right)-P\left(\frac{\sum_{j=1}^{k+L}Z_{j}}{\sqrt{\alpha N+k}}\leq\frac{k}{\alpha N+k}\right)\right]$
$>\prod_{k=\alpha N\delta_{N}}^{N}\left[1-O\left(\frac{1}{N}\right)\right]$
@@
$\log\prod_{k=1}^{\alpha N\delta_{N}}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+\alpha N}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)$
$=\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log P\left(\alpha_{k}^{(N)}\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)\leq\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{k+\alpha N}\left(Z_{j}-1\right)\leq\alpha_{k}^{(N)}\left(\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)\right)$
@@
where, for fixed $k\in\left[0,\alpha N\delta_{N}\right]$, the factor $\alpha_{k}^{(N)}=\sqrt{\frac{\alpha N}{\alpha N+k}}$ goes to one as $N$ goes to infinity.
@@
Furthermore, using the classical Edgeworth expansion, in line with equation $(9)$ in [22], this implies that
$\log P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)$
$=\log\left[\int_{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}^{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{\alpha N}}\sup_{|c|\leq H}(\phi_{N}^{(R)}(c))^{2}e^{-\frac{(\phi_{N}^{(R)}(c))^{2}}{2}}\right)$
$+O\left(\frac{1}{\alpha N}\right)+O\left(\phi_{N}^{(r)}(y)(\alpha N)^{-3/2}\right)+O\left(e^{-\frac{(\phi_{N}^{(r)}(y))^{2}}{2}}\right)$
@@
where, as in [22], for $x$ restricted as in $|x|\leq H$ for some large positive $H$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq H}\phi_{N}^{(R)}(x)$.
Then,
$\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log\left[\int_{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}^{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
$=\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log\left[1-\int_{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt-\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
In the limit as $N$ goes to infinity,
$\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log\left[1-\int_{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt-\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
$=\sum_{k=1}^{\sqrt{2\alpha N\log N}}\left[-\int_{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt-\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
$+\sum_{k=1}^{\sqrt{2\alpha N\log N}}O\left(\left(\int_{\phi_{N}^{(R)}(x)+C_{\alpha,N}+\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt+\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)^{2}\right)$
$=-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)+O\left(\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right)\right)$
$-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
$=-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)+O\left(\frac{e^{-\frac{\left(\phi_{N}^{(r)}(y)\right)^{2}}{2}}}{\phi_{N}^{(r)}(y)}\right)\right)$
$-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
This implies that,
$\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log\left[1-\int_{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt-\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
$=-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
$-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
where $\Phi_{N}^{(R)}(x)=\phi_{N}^{(R)}(x)+C_{\alpha,N}$.
The leading order sum as the Riemann integral of last equation is,
$-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
$-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
$=-\sqrt{\alpha N}\int_{\Phi_{N}^{(R)}(x)}^{\Phi_{N}^{(R)}(x)-\sqrt{2\log(N)}}\left(\int_{u}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{u}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)\right)du$
$-\sqrt{\alpha N}\int_{\phi_{N}^{(r)}(y)}^{\phi_{N}^{(r)}(y)+\sqrt{2\log(N)}}\left(\int_{-\infty}^{-u}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-u}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)\right)du+E_{N}$
@@
$E_{N}=O\left(\left(\sqrt{\frac{\log N}{\alpha N}}\right)\vee\left(\sqrt{\log{N}}\sup_{|c|\leq H}(\Phi_{N}^{(R)}(c))^{2}e^{-\frac{(\Phi_{N}^{(R)}(c))^{2}}{2}}\right)\right)$
Remark. For $\alpha=N^{m},(\rho^{2}-1)\sqrt{\alpha N}=\frac{1}{\alpha}\sqrt{\alpha N}=\frac{1}{N^{m-\frac{m}{2}-\frac{1}{2}}}=\frac{1}{N^{\frac{m-1}{2}}}$. In the limit as $N$ goes to infinity, with $\alpha>0$, under the condition that
    $\exists m_{1}>1$, $\forall m>m_{1}$ and $\alpha=O\left(N^{m}\right)$, as $N$ goes to infinity, $\phi_{N}^{(R)}(x)=o\left(\sqrt{\alpha N}\right)>\sqrt{2\log(N)}$.
This assumption is used such that the term $C_{\alpha,N}$, goes to zero as $N$ goes to infinity.
Following the framework presented in [22],
$\log F_{X}^{(\infty)}(x)$ $=-\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(R)}(x)}^{+\infty}\left(\int_{u}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{u}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)\right)du$
$=-\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(R)}(x)}^{+\infty}\left(\int_{u}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)du\times\left(1+O\left(\frac{1}{\left(\phi_{N}^{(R)}(x)\right)^{2}}\right)\right)$
$=-\lim_{N\rightarrow+\infty}\sqrt{\frac{\alpha N}{2\pi}}\frac{e^{-\frac{\left(\phi_{N}^{(R)}(x)\right)^{2}}{2}}}{\left(\phi_{N}^{(R)}(x)\right)^{2}}\times\left(1+O\left(\frac{1}{\left(\phi_{N}^{(R)}(x)\right)^{2}}\right)\right)$
$=-\exp(-x)$
with, $x\in\mathbb{R}$,
$f_{N}^{2}(x)=\frac{1}{2\rho^{2}}\left(\log\left(\frac{e^{x}\sqrt{N/(2\pi)}}{\log(N)}\right)\right)$
And,
$\log\left(1-F_{Y}^{(\infty)}(y)\right)$ $=-\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(r)}(y)}^{+\infty}\left(\int_{-\infty}^{-u}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-u}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)\right)du$
$=-\exp(y)$
with, $y\in\mathbb{R}$,
$f_{N}^{2}(y)=\frac{1}{2}\left(\log\left(\frac{e^{-y}\sqrt{N/(2\pi)}}{\log(N)}\right)\right)$
@@
$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\rho^{2}\alpha N}}+\frac{x}{2\sqrt{2\rho^{2}\alpha N\gamma_{\alpha,N}}}\right]$
@@
This proves the independence of the scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$ of a $N\times N$ matrices from the complex induced Ginibre ensemble with proportional rectangularity index as $N$ goes to infinity.